OpenStudy (anonymous):

Help @ModernPraise97 This is my problem a^2-25/a+5=6/3 solve by cross multiplying

OpenStudy (anonymous):

This time I got 7 but it still doesn't check.

OpenStudy (anonymous):

ok show me your work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for some reason it won't let me draw so give me one sec and i think it actually does check:)

OpenStudy (anonymous):

it wont let me draw or do equations... I think its my stupid computer... i hate it lol

OpenStudy (anonymous):

a^2-25=2(a+5) (a+5)(a-5)=2(a+5) a-5=2 a=7 Check: 7^2-25/7+5=6/3 49-25/12=6/3 24/12=6/3 2=2 lol

OpenStudy (anonymous):

It is kind of funny how people type lol all the time and they probably aren't really laughing out loud. HA HA

OpenStudy (anonymous):

TRUE haha

OpenStudy (anonymous):

let me guess u weren't really laughing HA HA HA

OpenStudy (anonymous):

are u still available or no???

OpenStudy (anonymous):

hey sorry ummm im giong to bed in about 15 min is it quick?

OpenStudy (anonymous):

Oh that ok hit the hay:) just one question though do u come on this sight often cause i would really like to get your help at other times???

OpenStudy (anonymous):

yeah man I just got my account a few hours ago but im starting the new semester next week so I will have alot of down time and after tommorrow I will have the rest of the days off

OpenStudy (anonymous):

for this week

OpenStudy (anonymous):

so you are in college or highschool? And I'm actually not a man:) my first name is Taelor:)lol

OpenStudy (anonymous):

So how would i reach you for help???

OpenStudy (anonymous):

ah ha! ok ill keep that in mind lol I am a junior in high school and do you have a phone? you could text me if you need help

OpenStudy (anonymous):

Ah, i could use my dad's or mom's but I don't have one:) What is the number??? And just wondering where are you from???

OpenStudy (anonymous):

850-273-4246 text only its a pinger number I am from Florida and still live in Florida :)

OpenStudy (anonymous):

Oh i wish i could be in florida right know i Am from northern Minnesota and the temp right know is 0 degrees brrr!!! :)

OpenStudy (anonymous):

By the way I am a sophomore:)

OpenStudy (anonymous):

ooooh my friend lives in Michigan and she says it was 5 when she went to take a walk... I want to be in the cold so bad... I have never seen snow in my life haha

OpenStudy (anonymous):

Are u serious??? I could not have Christmas without it!!! In the pic of you you look like a college dude:)

OpenStudy (anonymous):

just saying...

OpenStudy (anonymous):

Ive never had christmas with it! lol oh haha yeah people tell me Im older than I look...

OpenStudy (anonymous):

I had one of my friend's friend think i was My friends little sister even though I am actually older. I am 16 and I don't know what people think know,but i wouldn't be surprised if they guessed me at 14:) It gets annoying!!!

OpenStudy (anonymous):

awww! haha you said yoru name was Taelor? thats an awesome spelling!

OpenStudy (anonymous):

Wait people tell you your older then u look? don't you mean the other way around? Oh thanks! That is a little annoying though when I go on trips though because i can't get sooviners(how ever u spell that:) My other nick name is Bud:)

OpenStudy (anonymous):

I guess I am little bit of A tom boy but other times I'm pretty girly:)

OpenStudy (anonymous):

oh yeah the other way around sorry lol brain friend tonight

OpenStudy (anonymous):

*fried

OpenStudy (anonymous):

He he I love fishing that is what i have done since I was probably 5. I love the outdoors and speaking of that I am goig skiing on friday. Please tell me you have been skiing:)

OpenStudy (anonymous):

Snow skiing that is:)

OpenStudy (anonymous):

nope! never seen snow remember?

OpenStudy (anonymous):

Man alive! U r missing out! Oh you just have to see snow soon. And go skiing:) Do you have any siblings? I have 2 bros and 1 sis. all married and gone so pretty lonely around here:)

OpenStudy (anonymous):

going to Atlanta next month maybe ill be lucky enough to see some snow. I have 2 bros and 1 sis too!!!! im the oldest though..

OpenStudy (anonymous):

oh great ok I have to go to bed... I have exams in the morning and moms up my back :/

OpenStudy (anonymous):

Anyway i am losing my brain talking back and forth. I'm suppose to be doing homework. It was great talking to u, beach boy:) and I'll text you if i need your help . Taelor

OpenStudy (anonymous):

alrighty snow white... nice talking to you too :) bye

OpenStudy (anonymous):

Hey thanks snow white I like that beach boy:)

OpenStudy (anonymous):

Good night:)

OpenStudy (anonymous):

Hey what is your name???